OpenStudy (anonymous):

Simplify: (-2x^2y^3z)^3=

OpenStudy (anonymous):

When you raise powers to a power, you multiply the exponents: \[\huge (ab^kc^n)^r = a^rb^{k\cdot r}c^{n\cdot r}\]

OpenStudy (anonymous):

i * them o ok thanks

OpenStudy (anonymous):

Easy to forget the coefficient on this one...so = -8x^6 * y^9 * z^3